(4+x^2)dy +4dx=0

1
Expert's answer
2021-11-14T18:16:24-0500
"dy=-\dfrac{4}{4+x^2}dx"

Integrate


"\int dy=-\int\dfrac{4}{4+x^2}dx"

"\int\dfrac{4}{4+x^2}dx=\int\dfrac{1}{1+(x/2)^2}dx=2\arctan (x/2)+C"

Then


"y=-2\arctan(x/2)+C"